A project has the following cash flows. What is the payback period?

Year   Cash flow

0   -14500

1 2200

2   4800

3   6500

4 7500

My question is that should I find positive value in any year when I add future value from the initial value (Year 0) to make a positive value ? Could anyone explain this question?
Answer #1

As per the question, we need to calculate simple payback period, in case of discounted payback period, we need to discount the future values and then calculate the payback period. In this case, we will follow the following process to calculate the simple payback period.

Net cash flow:
Year 0: -14500
Year 1: 2200
Year 2: 4800
Year 3: 6500
Year 4: 7500

Cumulative net cash flows:
Year 0: -14500
Year 1: 2200-14500=-12300
Year 2: 4800-12300=-7500
Year 3: 6500-7500=-1000
Year 4: 7500-1000=6500

Payback period= Full time until recovery + (Unrecovered cost at the beginning of last year)/(Cash flow during the last year)

Full time until recovery=3
Unrecovered cost at the beginning of last year=1000
Cash flow during the last year=7500

Payback period=3+1000/7500=3.133 years